OpenStudy (anonymous):

What is the solution set for |2x – 8| 10?

OpenStudy (anonymous):

> or equal to 10

jimthompson5910 (jim_thompson5910):

\[\large |2x – 8|\ge 10\] \[\large 2x – 8\ge 10 \text{ or } 2x – 8\le -10\] \[\large 2x \ge 10+8 \text{ or } 2x \le -10+8\] \[\large 2x \ge 18 \text{ or } 2x \le -2\] \[\large x \ge \frac{18}{2} \text{ or } x \le -\frac{2}{2}\] \[\large x \ge 9 \text{ or } x \le -1\]